HETAV PATEL

Grade 11,

A and B are two points .The position vector of A is 6b-2a.A point P divides the line AB in the ratio 1:2 If a-b is the p.v. of P then the P.V. of B is given by

A and B are two points .The position vector of A is 6b-2a.A point P divides the line AB in the ratio 1:2 If a-b is the p.v. of P then the P.V. of B is given by

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear palllavi

use section formula.

let position vector of B is r

so position vector of P = (1* position vector of B  + 2 * position vector of a)/(2+1)

     a-b  = (r +2(6b-2a))/2

 2a-2b  =r +12b -4a

 r= 6a-14b

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin